# Cox reduction and confidence sets of models: a theoretical elucidation

### v2

@@ 9.2 Cox reduction with unknown variance
$T_{Y:K}$ $=$ $\hat{\sigma}_{\mathit{K}}^{-1}D_{\mathit{K}}^{-1/2}\hat{\theta}_{Y:\mathit{K}},$
$\hat{\sigma}^{2}_{\mathit{K}}$ $=$ $\frac{\|Y-X_{\mathit{K}}\hat{\theta}_{Y:\mathit{K}}\|_{2}^{2}}{n-|\mathit{K}|}$
@@ 9.5 Prediction
The argument for confidence sets of models is somewhat weakened when prediction is the primary goal. However, stability of the predictor over time and in different contexts is important and more likely with a model that has an underlying interpretation as well as immediate predictive success. By consideration of prediction intervals for each model in $\mathcal{M}$, we account both for model uncertainty and statistical uncertainty. Overlapping prediction intervals are reassuring, while those differing considerably point to instability of the predictor under alternative circumstances. See Section S3 of the supplementary material for an example.

### v3

@@ 9.2 Cox reduction with unknown variance
$T_{Y:K}=\hat{\sigma}_{\mathit{K}}^{-1}D_{\mathit{K}}^{-1/2}\hat{\theta}_{Y:\mathit{K}},\quad\quad\hat{\sigma}^{2}_{\mathit{K}}=\frac{\|Y-X_{\mathit{K}}\hat{\theta}_{Y:\mathit{K}}\|_{2}^{2}}{n-|\mathit{K}|}$
@@ 9.5 Prediction
The argument for confidence sets of models is somewhat weakened when prediction is the primary goal. However, stability of the predictor over time and in different contexts is important and more likely with a model that has an underlying interpretation as well as immediate predictive success. By consideration of prediction intervals for each model in $\mathcal{M}$, we account both for model uncertainty and statistical uncertainty. Overlapping prediction intervals are reassuring, while those differing considerably point to instability of the predictor under alternative circumstances. See Section S3 of the supplementary material [36] for an example.
@@
Supplementary Material. The supplement provides pseudo-code for the procedure, simulation results and proofs of key results.